OpenStudy (anonymous):

How do I solve for w in 13+w/7=-18? w/7 is a fraction

OpenStudy (anonymous):

Do you know how to get started?

OpenStudy (nurali):

\[13+\frac{ w }{ 7 }=-18\]

OpenStudy (nurali):

\[\frac{ 13*7+w }{ 7 }=-18\]

OpenStudy (nurali):

\[91+w=-126\]

OpenStudy (nurali):

\[w=-126-91\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."